# Distinguishing Predictive and Generative AI in Regulation

### v1

@@ 4. Recommendations
AI regulation that relies on technical proxies risks not just becoming outdated but being both under- and over-inclusive (Bommasani, 2023). Regulation is underinclusive when it leaves out systems that achieve similar concerning outcomes through alternative methods (Crootof and Ard, 2021). For instance, relying on compute thresholds to identify frontier AI models may inadvertently exclude systems that cause harm through less resource-intensive methods (e.g., misinformation bots) (Newsom, 2024). Regulation is overinclusive when it captures more systems than intended. For example, the same compute thresholds could sweep in benign applications like climate modeling that happen to share high computational demands. The challenge in defining the regulatory scope is compounded by the broad range of contexts in which the same AI systems can be deployed to serve vastly different purposes (Harris, 2021). The generality and adaptability of generative AI therefore underscore the need for regulatory frameworks that account for the specific contexts and outcomes of their use (Murdick et al., 2020).Tech-neutral regulation also helps address the challenge of regulatory arbitrage, where actors might switch to alternative technical approaches to circumvent tech-specific rules while achieving the same concerning outcomes (Crootof and Ard, 2021; Burk, 2016). This is especially relevant given the rapid pace of AI development, where new architectures and methodologies emerge frequently. Tying regulations to specific technologies creates incentives for companies to design their systems in ways that fall outside the regulatory scope. One example that highlights the limits of tech-specific regulation is American Broadcasting Companies v. Aereo (ABC v. Aereo, Inc., 573 U.S. 431, 2014). Aereo was a company that designed a new technology to transmit broadcast television signals while technically complying with a legal precedent set in Cablevision. In Cablevision, the court ruled that a content recording service did not violate copyright law if a consumer initiated the recording and the system created a unique copy for each user that was accessible only to them (20th Century Fox v. Cablevision, 536 F.3d 121, 2008). Based on this ruling, Aereo assigned each user their own miniature antenna and made a separate copy of broadcast content for each viewer. This allowed mass transmission, as thousands of antennas could be tuned to the same signal simultaneously, yet each transmission remained technically private and user-initiated. Its system was effectively a technical workaround designed to achieve the same outcome as a cable service but without being subject to the same legal obligations. The Supreme Court ultimately held that Aereo’s service did violate copyright law, concluding that its operations were functionally equivalent to those of cable providers and therefore fell within the scope of the Copyright Act.

### v2

@@ 4. Recommendations
AI regulation that relies on technical proxies risks not just becoming outdated but being both under- and over-inclusive (Crootof and Ard, 2021). [^5] Considered more precisely, this is a point about legal rules and standards, not tech neutrality (Crootof and Ard, 2021, p. 406). Where regulations are more rule-like, they can be over- and under-inclusive. We make the point here about tech-specificity because we think it is likely that tech-specificity leads to rules more often than standards. This is both because the tendency for tech-specific regulation is already motivated by a desire for easier adminsitrability and because technical standards and specifications are often stated in quantitative terms. Examples of tech-specific standard-based regulation may become more prevalent if governments, for example, mandate that LLMs have output filters that accomplish some goal like fairness or a more complicated set of goals in the vein of Anthropic’s attempt at “Constitutional AI”, (Bai et al., 2022) but for now, we think the assumption holds. Regulation is underinclusive when it leaves out systems that achieve similar concerning outcomes through alternative methods. For instance, relying on compute thresholds to identify frontier AI models may inadvertently exclude systems that cause harm through less resource-intensive methods (e.g., misinformation bots) (Newsom, 2024). Regulation is overinclusive when it captures more systems than intended. For example, the same compute thresholds could sweep in benign applications like climate modeling that happen to share high computational demands. The challenge in defining the regulatory scope is compounded by the broad range of contexts in which the same AI systems can be deployed to serve vastly different purposes (Harris, 2021). The generality and adaptability of generative AI therefore underscore the need for regulatory frameworks that account for the specific contexts and outcomes of their use (Murdick et al., 2020).
Tech-neutral regulation also helps address the challenge of regulatory arbitrage, where actors might switch to alternative technical approaches to circumvent tech-specific rules while achieving the same concerning outcomes (Crootof and Ard, 2021; Burk, 2016). This is especially relevant given the rapid pace of AI development, where new architectures and methodologies emerge frequently. Tying regulations to specific technologies creates incentives for companies to design their systems in ways that fall outside the regulatory scope (Yew et al., 2025).
One example that highlights the limits of tech-specific regulation is the European Union’s legislation on genetically modified organisms (GMO) (Parliament and the Council of the European Union, 2001). The legislation targets crops that have been altered by recombinant DNA technology to address public concerns over potential health risks. In response, seed producers have turned away from recombinant DNA technology toward mutagenesis, using mutagenic chemicals or radiation treatments to induce random genetic changes in plants (Kaskey, 2013a). While these methods are not considered recombinant DNA technology, they achieve the same outcome of producing genetically modified crops with commercially desirable traits. Ironically, mutagenic crops may pose comparable or even greater risks than GMO crops on food safety and the environment (Kaskey, 2013b; Morris, 2007). This case demonstrates how tech-specific regulation can incentivize the development of what Dan Burk calls “perverse innovations” that “ingeniously dodge the intended outcome of regulation, while still formally adhering to the text of the regulation (Burk, 2016).”